# Multi-Token Joint Speculative Decoding for Accelerating Large Language Model Inference

### v1

@@ 5.1 Comparison with Baselines
Table 2 shows the main results of our experiments. The standard deviations in the table are computed by repeating the experiment four times. First, we compare the output perplexity of different algorithms. We can see that the perplexity of greedy decoding, vanilla speculative decoding, and Spectr decoding are close because they have identical sampling distributions for each token. Meanwhile, BiLD approximates the sampling distribution of greedy decoding but yields better perplexity. It is because we set a strict acceptance threshold for BiLD decoding, which lowers the acceptance rate but ensures every token has a high probability in the large model. It can be viewed as having an additional warping operation towards the sampling distribution. Nevertheless, there is a significant gap between the perplexity of MJSD and other baselines. On average, the output perplexity of MJSD is $21.2\%$ lower than that of BiLD. It provides strong evidence that MJSD has better effectiveness than existing speculative decoding methods that aim to have identical sampling distribution as greedy decoding.
Next, we compare the speeds of different algorithms. While the speeds of all speculative decoding baselines are close, MJSD is significantly faster. On average, MJSD is 2.21×2.21$\times$2.21 × faster than greedy decoding, 1.49×1.49$\times$1.49 × faster than vanilla speculative decoding, 1.55×1.$5$ faster than BiLD, and 1.64×1.64$\times$1.64 × faster than Spectr. The speed improvement is because MJSD has a larger average acceptance length at each iteration, as MJSD accepts the longest sub-sequence that passes verification. Figure 3 shows the average acceptance length of vanilla speculative decoding and MJSD. We can see that the acceptance length of MJSD is on average 2.15×2.15$\times$2.15 × larger than that of vanilla speculative decoding.
Finally, we focus on energy efficiency. We can see that all speculative decoding algorithms have significantly better energy efficiency over greedy decoding. It supports our analysis that speculative decoding can improve the LLM inference energy efficiency by reducing the number of memory operations. In addition, we can see that the energy efficiency of MJSD is significantly better than that of all baseline algorithms. We believe it is also because of the longer acceptance sequence at each iteration. The number of memory operations is proportional to the number of LLM runs, which is inversely proportional to the average acceptance length. Since by always accepting the longest prefix that passes verification, the energy consumption of MJSD is on average 1.62×1.62$\times$1.62 × smaller than that of vanilla speculative decoding, and 2.84×2.84$\times$2.84 × smaller than that of greedy decoding.
Figure 3: Average number of accepted tokens per iteration of speculative decoding and MJSD. Each iteration generates 4 draft tokens.
@@ 5.2 Ablation Study
5.2.1 Effects of Number of Beams
5.2.2 Effects of Acceptance Threshold
Figure 4: Effect of acceptance threshold on output perplexity and decoding speed on ChatGPT-Prompts dataset.
Next, we evaluate the effect of acceptance threshold $\tau$. Intuitively, when we increase $\tau$ from 0 to 1, the acceptance criterion becomes more strict, the efficiency drops while the output perplexity increases. Surprisingly, this expectation is only partially correct. As shown in Figure 4, the efficiency indeed drops when $\tau$ increases. However, the perplexity increases when $\tau$ is close to 1. When $\tau=1$, all the draft tokens are rejected, which makes MJSD equivalent to greedy decoding. Similarly, when $\tau$ is close to 1, the advantage of multi-token joint greedy decoding on effectiveness disappears, hence the perplexity becomes close to the perplexity of greedy decoding. Another surprising observation is that the perplexity of MJSD is good when $\tau=0$. When $\tau=0$, MJSD is equivalent to generating $\gamma$ tokens using beam decoding with the small model, then generating an additional token with the large model. The fact that MJSD achieves good perplexity when $\tau=0$ can be explained by the fact that “even when a small model is an order of magnitude smaller than a large model, only a small fraction of the small model’s predictions deviate from those of the large model” [17, 20]. Moreover, when $\tau$ ranges from $0.1$ to $0.9$, the performance of MJSD is relatively stable, suggesting that MJSD is not sensitive to the acceptance threshold.
@@ 6 Related Work
Efficient Decoding Inference. There are extensive studies on improving large model inference efficiency. Well-known methods include model quantization [9, 22], model pruning [10, 28], and model distillation [15]. Despite achieving significant speed-ups, a common drawback of these methods is that they have to sacrifice the model’s effectiveness.
A direction closer to our work is non-autoregressive decoding. It is first proposed by [12] to generate multiple tokens in parallel. That is, the model simultaneously predicts $p(x_{t+k}|x_{1:t})$ ($k=1,2,\ldots$). Subsequent studies further improved the performance of parallel decoding by incorporating additional information [35, 31, 21] or using additional iterations to refine predictions [11, 19, 13]. However, these works require continuous training of the model and usually either compromise the model effectiveness or require task-dependent techniques to achieve comparable performance [17].
Speculative Decoding. Speculative decoding was recently proposed in [20, 5] as a way to accelerate LLM inference. Spectr [32] enhances speculative decoding by letting the small model generate multiple i.i.d. draft sequences. While speculative decoding and Spectr use the large model to verify all the tokens drafted by the small model, BiLD [17] only calls the large model when the probability output by the small model is below a pre-defined threshold $\tau_{1}$. The large model rejects a token if its negative log-likelihood is larger than threshold $\tau_{2}$. SpecInfer [24] uses one or multiple small models to generate a draft token tree to increase the average acceptance length for each iteration. All these methods can be perceived as exact or approximate versions of sampling tokens from the conditional distribution $p(x_{t}|x_{<t})$. Therefore, their output perplexity is bounded by greedy decoding.
An orthogonal direction to improve speculative decoding is to improve the effectiveness of the small draft model. It is obvious that if more draft tokens are accepted, the overall inference speed will increase. BiLD [17] uses a model prediction alignment technique to better train the small model. Liu et al. [23] propose online speculative decoding to continually update the draft model based on observed input data. Instead, Rest [14] uses a retrieval model to produce draft tokens. The improvement of small models can benefit all speculative decoding algorithms, including our method.
@@ 7 Conclusion
We introduce multi-token joint speculative decoding that significantly enhances output quality with better time and energy efficiency. A distinctive aspect of our work is the exploration of speculative decoding’s impact on inference energy consumption, an often neglected area in existing studies. Experiment results demonstrate our method achieves significant energy reduction. This research contributes not only a novel decoding approach but also valuable insights for optimizing LLM deployment in real-world applications where considerations of both quality and efficiency are crucial.
8 Limitations and Impact Statements
Limitations. This paper mainly considers improving the output perplexity of decoding algorithms. Under the assumption that the model is well-trained for downstream tasks, improving output perplexity usually leads to better downstream effectiveness. But if the model is not well-trained, improving the perplexity may not necessarily improve the downstream effectiveness. However, it is mainly the problem of the model itself. In real-world applications, it is reasonable to assume the model is properly trained for downstream tasks.
Impact Statements. The goal of this work is to advance the field of Large Language Model (LLM) Inference, which has received tremendous attention from both academia and industry. However, LLMs also have received many critiques, including their extremely large carbon footage emission during both training and inference. Our work pays special attention to their energy consumption during inference to provide high-quality and fast-inference LLMs with reduced energy consumption, which has become a serious concern with the rapid increase of LLM-related workloads in the past few years.
Appendix A Pseudocode of MJSD

### v2

@@ 5.1 Comparison with Baselines
Table 4 shows the primary results of our experiments while Table 5 details the block efficiency, defined as the average number of tokens generated per iteration, of different methods. The standard deviations in the tables are computed by repeating each experiment four times [^4] For MT-Bench, the standard deviation also accounts for variations across different tasks.. First, we observe that MTAD is significantly more efficient than all baselines in terms of both energy and time. On one hand, the energy consumption of MTAD is on average 1.54×1.54$\times$1.54 × smaller than that of vanilla speculative decoding. On the other hand, MTAD is 1.101.71×1.10-1.71$\times$1.10 - 1.71 × faster than vanilla speculative decoding, 1.38×1.38$\times$1.38 × faster than BiLD, 1.59×1.59$\times$1.59 × faster than Spectr, and 1.60×1.60$\times$1.60 × faster than SpecInfer. SpecInfer and Spectr has better block efficiency than vanilla speculative decoding, but are slower. This may be due to the fact that they have to verify multiple draft sequences, which introduces extra overhead and may not be perfectly parallelized, especially when the memory overhead exceeds the GPU memory capacity. Meanwhile, MTAD has the best block efficiency without causing any extra overhead, hence it is significantly more efficient.
Table 5: Average number of tokens generated at each iteration across all datasets.
Llama-2 OPT
spec 2.02$\pm$0.05 2.60$\pm$0.06
BiLD 1.83$\pm$0.10 2.68$\pm$0.36
Spectr 2.73$\pm$0.43 3.45$\pm$0.42
SpecInfer 2.74$\pm$0.46 3.45$\pm$0.40
MTAD 3.17$\pm$0.43 4.30$\pm$0.03
@@
Next, we compare the output perplexity of different algorithms. The perplexity scores of vanilla speculative decoding, SpecInfer, and Spectr are close since their sampling distributions are equivalent. Meanwhile, BiLD approximates the sampling distribution of single-token multinomial sampling but yields better perplexity. It is because we set a strict acceptance threshold for BiLD, which lowers the acceptance rate but ensures every token has a high probability in the large model, thus improving the overall likelihood. More importantly, there is a significant gap between MTAD and other baselines. On average, the perplexity of MTAD is $21.2\%$ lower than that of speculative decoding.
Table 6: Downstream task scores of speculative decoding and MTAD. All the scores are higher the better.
spec MTAD
CD Rouge-L 0.114 0.118
SP EA 11.5 13.0
MT Humanities 2.95 3.15
Extraction 1.80 2.50
Roleplay 3.10 3.80
Math 1.10 1.00
Coding 1.25 1.10
Reasoning 3.80 3.15
STEM 2.85 3.10
Writing 3.80 3.65
Average 2.58 2.68
In addition, to show MTAD indeed improves the downstream effectiveness, we compare the performance metrics of speculative decoding and MTAD on CNNDM, Spider, and MT-Bench datasets. [^5] All speculative baselines are equivalent to multinomial sampling We exclude the other two datasets as they lack explicit downstream metrics. And we exclude the results of OPT models due to their consistently poor performance across all evaluated datasets. As illustrated in Table 6, MTAD outperforms speculative decoding across all three datasets, thereby validating our claim that MTAD achieves superior effectiveness compared to conventional decoding methods that rely on single-token distributions.
@@ 5.2.1 Number of Beams
5.2.2 Acceptance Threshold
Figure 3: Effect of acceptance threshold on output perplexity and decoding speed.
Next, we evaluate the effect of acceptance threshold $\tau$. Intuitively, when we increase $\tau$ from 0 to 1, the acceptance criterion becomes more strict, the efficiency drops while the output perplexity increases. Surprisingly, this expectation is only partially correct. As shown in Figure 3, the efficiency indeed drops when $\tau$ increases. However, the perplexity increases when $\tau$ is close to 1. When $\tau=1$, all the draft tokens are rejected, which makes MTAD equivalent to multinomial sampling. Similarly, when $\tau$ is close to 1, the advantage of multi-token joint decoding on effectiveness disappears, hence the perplexity becomes close to the perplexity of multinomial sampling. Another surprising observation is that the perplexity of MTAD is good when $\tau=0$. When $\tau=0$, MTAD is equivalent to generating $\gamma$ tokens using beam decoding with the small model, then generating an additional token with the large model. The fact that MTAD achieves good perplexity when $\tau=0$ can be explained by the fact that “even when a small model is an order of magnitude smaller than a large model, only a small fraction of the small model’s predictions deviate from those of the large model” (Kim et al., 2023; Leviathan et al., 2023). Moreover, when $\tau$ ranges from $0.1$ to $0.9$, the performance of MTAD is relatively stable, suggesting that MTAD is not sensitive to the acceptance threshold.
@@ 6 Related Work
Efficient Decoding Inference. There are extensive studies on improving large model inference efficiency. Well-known methods include model quantization (Frantar et al., 2022; Lin et al., 2023), model pruning (Gale et al., 2019; Sanh et al., 2020), and model distillation (Hinton et al., 2015). Despite achieving significant speed-ups, a common drawback of these methods is that they have to sacrifice the model’s effectiveness.
A direction closer to our work is non-autoregressive decoding. It is first proposed by (Gu et al., 2017) to generate multiple tokens in parallel. That is, the model simultaneously predicts $p(x_{t+k}|x_{1:t})$ ($k=1,2,\ldots$). Subsequent studies further improved the performance of parallel decoding by incorporating additional information (Wang et al., 2019; Sun et al., 2019; Li et al., 2019) or using additional iterations to refine predictions (Ghazvininejad et al., 2019; Lee et al., 2018; Guo et al., 2020). However, these works require continuous training of the model and usually either compromise the model effectiveness or require task-dependent techniques to achieve comparable performance (Kim et al., 2023).
Speculative Decoding. Speculative decoding was recently proposed in (Leviathan et al., 2023; Chen et al., 2023) as a way to accelerate LLM inference. Spectr (Sun et al., 2023) enhances speculative decoding by letting the small model generate multiple i.i.d. draft sequences. While speculative decoding and Spectr use the large model to verify all the tokens drafted by the small model, BiLD (Kim et al., 2023) only calls the large model when the probability output by the small model is below a pre-defined threshold $\tau_{1}$. The large model rejects a token if its negative log-likelihood is larger than threshold $\tau_{2}$. SpecInfer (Miao et al., 2023) uses one or multiple small models to generate a draft token tree to increase the average acceptance length for each iteration. All these methods can be perceived as exact or approximate versions of sampling tokens from the conditional distribution $p(x_{t}|x_{<t})$. Therefore, their output perplexity is bounded by greedy decoding.
An orthogonal direction to improve speculative decoding is to improve the effectiveness of the small draft model. It is obvious that if more draft tokens are accepted, the overall inference speed will increase. BiLD (Kim et al., 2023) uses a model prediction alignment technique to better train the small model. Liu et al. (Liu et al., 2023) propose online speculative decoding to continually update the draft model based on observed input data. Instead, Rest (He et al., 2023) uses a retrieval model to produce draft tokens. An alternative way is to train additional heads in the large model to predict future tokens. Representative works include EAGLE (Li et al., 2024) and MEDUSA (Cai et al., 2024). Importantly, these works are orthogonal to speculative decoding techniques, including our proposed method. This orthogonality means that the improvements offered by more accurate draft tokens could be combined with our method for better effectiveness.
@@ 7 Conclusion
We introduce multi-token assisted decoding that significantly enhances output quality along with better time and energy efficiency. A distinctive aspect of our work is the exploration of speculative decoding’s impact on inference energy consumption, an often neglected area in existing studies. This research contributes not only a novel decoding approach but also valuable insights for optimizing LLM deployment in real-world applications where considerations of both quality and efficiency are crucial.
Appendix A Proof
A.1 Proof of Theorem 3.2
Proof.
$PPL(x_{1:\Gamma_{N}})$ $=\exp\left(-\frac{1}{\Gamma_{N}}\sum_{i=1}^{\Gamma_{N}}\log p(x_{i}|x_{1:i-1})\right)$ (10)
$=\exp\left(-\frac{N}{\Gamma_{N}}\frac{1}{N}\sum_{i=1}^{N}\log p(x_{\Gamma_{i-1}:\Gamma_{i}}|x_{1:\Gamma_{i-1}})\right)$
When $N\rightarrow\infty$, $\frac{\Gamma_{N}}{N}\rightarrow\bar{\gamma}$, and $\frac{1}{N}\sum_{i=1}^{N}\log p(x_{\Gamma_{i-1}:\Gamma_{i}}|x_{1:\Gamma_{i-1}})\rightarrow\mathbb{E}_{x_{1:t}\in\mathcal{X}}\sum_{\gamma}\sum_{x_{t+1:t+\gamma}}P(\gamma)\tilde{p}(x_{t+1:t+\gamma}|x_{1:t})\log p(x_{t+1:t+\gamma}|x_{1:t})=\mathbb{E}_{\gamma}L_{p}(\gamma,\tilde{p})$
A.2 Proof of Corollary 3.3
Proof.
For deterministic multi token sampling, $\tilde{p}_{multi}=\arg\max\circ p$, so we have
$L_{p}(\gamma,\tilde{p}_{multi})=E_{x_{1:t}\in\mathcal{X}}\max_{x_{t+1:t+\gamma}}\log p(x_{t+1:t+\gamma}|x_{1:t})$ (11)
Notice that deterministic greedy sampling can be seen as a special case of MJGD where $\tilde{p}_{single}(x_{t+1:t+\gamma}|x_{1:t})=1$ if and only if $x_{t+i}=\arg\max_{x}p(x|x_{1:t+i-1})$ for $i=1,\ldots,\gamma$. Let $x^{*}_{t+1:t+\gamma}$ be the tokens generated by deterministic MJGD and let $x^{\prime}_{t+1:t+\gamma}$ be the tokens generated by deterministic greedy decoding. For any fixed $\gamma$ and $x_{1:t}$, we have $\log p(x^{\prime}_{t+1:t+\gamma}|x_{1:t})\leq\max_{x_{t+1:t+\gamma}}\log p(x_{t+1:t+\gamma}|x_{1:t})=\log p(x^{*}_{t+1:t+\gamma}|x_{1:t})$. Therefore, $L_{p}(\gamma,\tilde{p}_{single})\leq L_{p}(\gamma,\tilde{p}_{multi})$. Then with Theorem 3.2, we know that the perplexity of greedy decoding will be higher. ∎
A.3 Proof of Lemma 3.4
We first prove the following Lemma.
Lemma A.1.
Let $PPL_{p}$ and $PPL_{q}$ denote the perplexity of tokens under distribution $p$ and $q$. When $N\rightarrow\infty$, we have
$\frac{PPL_{p}(x_{1:\Gamma_{N}})}{PPL_{q}(x_{1:\Gamma_{N}})}\leq\tau^{-\frac{1}{\bar{\gamma}}}$ (12)
where $\tau$ is the verification threshold.
Proof.
In the $i$-th iteration, the first $\gamma_{i}-1$ tokens are the accepted draft tokens and the last token is sampled from $p$. Based on our verification criteria, we know that for the accepted draft tokens, we have
$\frac{p(x_{\Gamma_{i-1}+1:\Gamma_{i-1}+\gamma_{i}-1}|x_{1:\Gamma_{i-1}})}{q(x_{\Gamma_{i-1}+1:\Gamma_{i-1}+\gamma_{i}-1}|x_{1:\Gamma_{i-1}})}\geq\tau.$ (13)
So,
$\frac{p({x_{1:\Gamma_{N}}})}{q({x_{1:\Gamma_{N}}})}\geq\tau^{N}\prod_{i=1}^{N}\frac{p(x_{\Gamma_{i}}|x_{1:\Gamma_{i}-1})}{q(x_{\Gamma_{i}}|x_{1:\Gamma_{i}-1})}$ (14)
Notice that
$\left(\prod_{i=1}^{N}\frac{p(x_{\Gamma_{i}}|x_{1:\Gamma_{i}-1})}{q(x_{\Gamma_{i}}|x_{1:\Gamma_{i}-1})}\right)^{\frac{1}{N}}=\exp\left(\frac{1}{N}\sum_{i=1}^{N}\log\left(\frac{p(x_{\Gamma_{i}}|x_{1:\Gamma_{i}-1})}{q(x_{\Gamma_{i}}|x_{1:\Gamma_{i}-1})}\right)\right)$ (15)
When $N\rightarrow\infty$, since the last token at each iteration is sampled from $p$, we have
$\frac{1}{N}\sum_{i=1}^{N}\log\left(\frac{p(x_{\Gamma_{i}}|x_{1:\Gamma_{i}-1})}{q(x_{\Gamma_{i}}|x_{1:\Gamma_{i}-1})}\right)\rightarrow\mathbb{E}_{p}\log\left(\frac{p(x_{\Gamma_{i}}|x_{1:\Gamma_{i}-1})}{q(x_{\Gamma_{i}}|x_{1:\Gamma_{i}-1})}\right)=KL(p,q)\geq 0$ (16)
So
$\left(\prod_{i=1}^{N}\frac{p(x_{\Gamma_{i}}|x_{1:\Gamma_{i}-1})}{q(x_{\Gamma_{i}}|x_{1:\Gamma_{i}-1})}\right)^{\frac{1}{N}}\geq 1$ (17)
Therefore,
$\frac{p({x_{1:\Gamma_{N}}})}{q({x_{1:\Gamma_{N}}})}\geq\tau^{N}$ (18)
Thus,
$\frac{PPL_{p}(x_{1:\Gamma_{N}})}{PPL_{q}(x_{1:\Gamma_{N}})}=\left(\frac{p({x_{1:\Gamma_{N}}})}{q({x_{1:\Gamma_{N}}})}\right)^{-\frac{1}{\Gamma_{N}}}\leq\tau^{-\frac{N}{\Gamma_{N}}}\rightarrow\tau^{-\frac{1}{\bar{\gamma}}}$ (19)
Now, we prove Lemma 3.4.
Proof.
$-\log PPL_{q}(x_{1:\Gamma_{N}})=\frac{1}{\Gamma_{N}}\sum_{i=1}^{N}(\log q(x_{\Gamma_{i-1}+1:\Gamma_{i}-1}|x_{1:\Gamma_{i-1}})+\log q(x_{\Gamma_{i}}|x_{1:\Gamma_{i}-1}))\$ (20)
When $N\rightarrow\infty$, since the first $\gamma_{i}-1$ tokens are sampled with beam decoding, we have
$\frac{1}{N}\sum_{i=1}^{N}\log q(x_{\Gamma_{i-1}+1:\Gamma_{i}-1}|x_{1:\Gamma_{i-1}}))$ $\rightarrow\mathbb{E}_{\gamma}\mathbb{E}_{x_{1:t\in\mathcal{X}}}\log q(x_{t+1:t+\gamma-1|x_{1:t}})$ (21)
$\geq(1-\epsilon)\mathbb{E}_{\gamma}\mathbb{E}_{x_{1:\Gamma_{i-1}}\in\mathcal{X}}\max_{x_{\Gamma_{i-1}+1:\Gamma_{i}-1}}q(x_{\Gamma_{i-1}+1:\Gamma_{i}-1}|x_{1:\Gamma_{i-1}}))$
$=(1-\epsilon)\mathbb{E}_{\gamma}L_{q}(\gamma-1,\arg\max\circ q)$
Since the last token at each iteration is sampled from $p$, we have
$\frac{1}{N}\sum_{i=1}^{N}\log q(x_{\Gamma_{i}}|x_{1:\Gamma_{i}-1}))\rightarrow\mathbb{E}_{x_{1:t}\in\mathcal{X}}\mathbb{E}_{p}\log q(x_{t+1}|x_{1:t})=-H(p,q)$ (22)
So
$-\log PPL_{q}(x_{1:\Gamma_{N}})\geq\frac{1-\epsilon}{\bar{\gamma}}\mathbb{E}_{\gamma,x_{1:\Gamma_{i-1}}\in\mathcal{X}}\max_{x_{t+1:t+\gamma}}q(x_{t+1:t+\gamma}|x_{1:t}))-\frac{H(p,q)}{\bar{\gamma}}$ (23)
$PPL_{q}(x_{1:\Gamma_{N}})\leq\exp\left(\frac{H(p,q)}{\bar{\gamma}}-\frac{1-\epsilon}{\bar{\gamma}}\mathbb{E}_{\gamma}L_{q}(\gamma-1,\arg\max\circ q)\right)$ (24)
A.4 Proof of Theorem 3.5
Proof.
We have
[…]
$=\tau^{-\frac{1}{\bar{\gamma}}}\exp\left(\frac{H(p,q)}{\bar{\gamma}}-\frac{1-\epsilon}{\bar{\gamma}}\mathbb{E}_{\gamma}L_{q}(\gamma-1,\arg\max\circ q)+\frac{1}{\bar{\gamma}}\mathbb{E}_{\gamma}L_{p}(\gamma,\arg\max\circ p)\right)$
Notice that $L_{p}(\gamma,\arg\max\circ p)\geq L_{p}(\gamma+1,\arg\max\circ p)$ for any $\gamma$. This is because for any $x_{1:t}$, $\max_{x_{t+1:t+\gamma}}\log p(x_{t+1:t+\gamma}|x_{1:t})\geq\max_{x_{t+1:t+\gamma+1}}\left(\log p(x_{t+1:t+\gamma}|x_{1:t})+\log p(x_{t+\gamma+1}|x_{1:t+\gamma})\right)=\max_{x_{t+1:t+\gamma+1}}\log p(x_{t+1:t+\gamma+1}|x_{1:t})$.
So
$\lim_{N\rightarrow\infty}\frac{PPL_{p}(x_{1:\Gamma_{N}})}{PPL_{p}(x^{*}_{1:\Gamma_{N}})}$ (26)
$\leq$ $\tau^{-\frac{1}{\bar{\gamma}}}\exp\left(\frac{H(p,q)}{\bar{\gamma}}+\frac{\epsilon}{\bar{\gamma}}\mathbb{E}_{\gamma}L_{p}(\gamma,\arg\max\circ p)+\frac{1-\epsilon}{\bar{\gamma}}(\mathbb{E}_{\gamma}L_{p}(\gamma,\arg\max\circ p)-\mathbb{E}_{\gamma}L_{q}(\gamma,\arg\max\circ q))\right)$
Since $\epsilon\leq 0$, and $L_{p}(\gamma,\arg\max\circ p)$ is the maximum log-likelihood, which is larger than the expected log-likelihood (i.e., negative entropy), we have
$\frac{\epsilon}{\bar{\gamma}}\mathbb{E}_{\gamma}L_{p}(\gamma,\arg\max\circ p)$ (27)
$=$ $\frac{\epsilon}{\bar{\gamma}}{E}_{\gamma}\mathbb{E}_{x_{1:t}\in\mathcal{X}}\max_{x_{t+1:t+\gamma}}\log p(x_{t+1:t+\gamma}|x_{1:t})$
$\leq$ $\frac{\epsilon}{\bar{\gamma}}{E}_{\gamma}\mathbb{E}_{x_{1:t}\in\mathcal{X}}\sum_{x_{t+1:t+\gamma}}p(x_{t+1:t+\gamma}|x_{1:t})\log p(x_{t+1:t+\gamma}|x_{1:t})$
$=$ $-\epsilon H(p)$
In addition
$\mathbb{E}_{\gamma}L_{p}(\gamma,\arg\max\circ p)-\mathbb{E}_{\gamma}L_{q}(\gamma,\arg\max\circ q)$ (28)
$=$ $\mathbb{E}_{\gamma}(L_{p}(\gamma,\arg\max\circ p)-L_{q}(\gamma,\arg\max\circ q))$
$=$ $\mathbb{E}_{\gamma}\left(\mathbb{E}_{x_{1:t}\in\mathcal{X}}\max_{x_{t+1:t+\gamma}}\log p(x_{t+1:t+\gamma}|x_{1:t})-\mathbb{E}_{x_{1:t}\in\mathcal{X}}\max_{x_{t+1:t+\gamma}}\log q(x_{t+1:t+\gamma}|x_{1:t})\right)$
$=$ $\mathbb{E}_{\gamma}\mathbb{E}_{x_{1:t}\in\mathcal{X}}\left(\max_{x_{t+1:t+\gamma}}\log p(x_{t+1:t+\gamma}|x_{1:t})-\max_{x_{t+1:t+\gamma}}\log q(x_{t+1:t+\gamma}|x_{1:t})\right)$
$\leq\mathbb{E}_{\gamma}\mathbb{E}_{x_{1:t}\in\mathcal{X}}\max_{x_{t+1:t+\gamma}}\left(\log p(x_{t+1:t+\gamma}|x_{1:t})-\log q(x_{t+1:t+\gamma}|x_{1:t})\right)$
$=\mathbb{E}_{\gamma}\mathbb{E}_{x_{1:t}\in\mathcal{X}}\max_{x_{t+1:t+\gamma}}\left(\sum_{i=1}^{\gamma}\log p(x_{t+i}|x_{1:t+i-1})-\log q(x_{t+i}|x_{1:t+i-1})\right)$
$\leq\mathbb{E}_{\gamma}\mathbb{E}_{x_{1:t}\in\mathcal{X}}U\gamma\quad(because\ \|\log p(x|x_{1:t})-\log q(x|x_{1:t})\|_{\infty}\leq U)$
$=U\bar{\gamma}$
And $H(p,q)=H(p)+KL(p\|q)$.
$KL(p\|q)$ $=\mathbb{E}_{x_{1:t}\in\mathcal{X}}\sum_{x}p(x|x_{1:t})(\log p(x|x_{1:t})-\log p(x|x_{1:t}))$ (29)
$\leq\mathbb{E}_{x_{1:t}\in\mathcal{X}}\sum_{x}p(x|x_{1:t})U\leq U$
So $H(p,q)\leq H(p)+U$. Therefore,
$\lim_{N\rightarrow\infty}\frac{PPL_{p}(x_{1:\Gamma_{N}})}{PPL_{p}(x^{*}_{1:\Gamma_{N}})}\leq\tau^{-\frac{1}{\bar{\gamma}}}\exp\left(\frac{(1-\epsilon\bar{\gamma})H(p)+(1-\epsilon+\bar{\gamma})U}{\bar{\gamma}}\right)$ (30)
A.5 Proof of Theorem 3.6
Proof.
Recall that we accept $x_{t+1:t+j}$ if and only if $\log p(x_{t+1:t+j}|x_{1:t})-\log q(x_{t+1:t+j}|x_{1:t})\geq\log\tau$. Since $\|\log p(x|x_{1:t})-\log q(x|x_{1:t})\|_{\infty}\leq U$, we have
$\log p(x_{t+1:t+j}|x_{1:t})-\log q(x_{t+1:t+j}|x_{1:t})\geq-jU$ (31)
Therefore $x_{t+1:t+j}$ is always accepted if $j\leq\frac{|\log\tau|}{U}$. So $\bar{\gamma}\geq\frac{|\log\tau|}{U}$
Appendix B Pseudocode of MJSD